OpenStudy (anonymous):

help plz your parents have a credit card with a balance of 3,287.90 at an interest rate of 14.5% apr they pay 1,200.00 each month on the due date until the card is paid off how many motnhs does it take to pay off the card and what is the total amount paid including interest

OpenStudy (anonymous):

@ranga can you help again plz

OpenStudy (anonymous):

$3.324.16 , because to calculate the monthly interest charged you divide the APR by 12 to get 1.2%.

OpenStudy (ranga):

Since this balance will get paid off in about 3 months, it would be easier to calculate the interest and the balance month by month rather than use some big formula. APR = 14.5%. Monthly interest rate = 14.5/12 = 1.2083% = 0.012083 Balance = $3,287.90 Interest after 1 month will be: 3,287.9 * 0.012083 * 1 = $39.73 Total balance = principal + interest = $3,287.90 + $39.73 = $3,327.63 First month payment = $1,200 Balance = $3,327.63 - $1,200 = $2127.63 Do the same until the balance is 0.